(1 point) Two bonds, each with a face value of $11000, are redeemable at par in t- years and priced to yield y2-696. Bond i has a coupon rate c2-11.1% and sells for $14523.39. Bond 2 has coupon rate c2-3.7% and sells for $ What is the value of P? Answer: $
Answer #1

Goal seek setting

Goal Seek Set cell: To value: By changing cell: SDS26 14523.39 SDS11 OK Cancel

Formula sheet

A B C D E F G H I J K
2
3 First time t needs to be calculated using the data given for bond 1
4 then t is to be used to calculated the value of bond 2.
5
6 For Bond 1:
7 Time t can be found using the goal seek function of excel such that the value of bond 1 is equal to the price given.
8 Par value (F) 11000
9 Coupon rate 0.111
10 Yield 0.06
11 Time to maturity 8.11652374782718 Years
12 Price of Bond 14523.39
13 Interest is paid once a year i.e. annual.
14 Annual coupon (C) =D8*D9
15 Annual Period (n) =D11
16 YTM (i) =D10
17 Current Value of the bond can be calculated by finding the present value of cash flows of bonds.
18 Cash Flow of Bonds can be written as follows:
19 Period 0 1 2 3 4 =D15
20 Cash Flow of Bonds =$D$14 =$D$14 =$D$14 =$D$14 =$D$14 =$D$14+D8
21
22 Current Value of Bond =C*(P/A,i,n)+F*(P/F,i,n)
23 Where, C is coupon, F is par value of bond, i is market rate and n is total number of periods.
24
25 Current Value of Bond =C*(P/A,i,n)+F*(P/F,i,n)
26 =D14*PV(D16,D15,-1,0)+D8*(1/((1+D16)^D15)) =D14*PV(D16,D15,-1,0)+D8*(1/((1+D16)^D15))
27
28 Value of t =ROUND(D11,0) Years
29
30 Calculation of price of bond 2:
31
32 Par value (F) 11000
33 Coupon rate 0.037
34 Yield 0.06
35 Time to maturity =D28 Years
36
37 Interest is paid once a year i.e. annual.
38 Annual coupon (C) =D32*D33
39 Annual Period (n) =D35
40 YTM (i) =D34
41 Current Value of the bond can be calculated by finding the present value of cash flows of bonds.
42 Cash Flow of Bonds can be written as follows:
43 Period 0 1 2 3 4 =D39
44 Cash Flow of Bonds =$D$38 =$D$38 =$D$38 =$D$38 =$D$38 =$D$38+D32
45
46 Current Value of Bond =C*(P/A,i,n)+F*(P/F,i,n)
47 Where, C is coupon, F is par value of bond, i is market rate and n is total number of periods.
48
49 Current Value of Bond =C*(P/A,i,n)+F*(P/F,i,n)
50 =D38*PV(D40,D39,-1,0)+D32*(1/((1+D40)^D39)) =D38*PV(D40,D39,-1,0)+D32*(1/((1+D40)^D39))
51
52 Hence value of bond 2 is =D50
53